The perimeter of a rectangle is 50 inches. If the ratio of its length to its width is 3 to 2, find the dimensions of the rectangle.
Answer by rfer(16322)

2(3x)+2(2x)=50
6x+4x=50
10x=50
x=50/10
x=5
3x=15 in
2x=10 in